CH/Agen/92/03

Agenda

Charissa Group Meeting

To be held on: 11 June 1992 To be held at: Oxford

Distribution: WNC,NMC,PVD,BRF,GJG,RAH,JSL,WDMR,GT,DLW

Agenda

0.  Minutes of Meeting of 29 April 1992

Matters Arising

1.  Update on NSC/SERC matters  (DLW)
2.  Report on Detector Tests  (WNC)
3.  Progress report on scintillators  (NMC)
4.  Progress report on silicon detectors  (BRF)
5.  Progress report on gas detector assemblies  (PVD or deputy)
6.  Progress report on chamber and gas system  (DLW)
7.  Progress report on electronics and DA  (RAH)
8.  Report on hybrid detector tests  (PVD or deputy)
9.  Plans for next runs (C+Mg, high energy, Si+Si)  (WNC,DLW,BRF)
10.  Update on Recent Analysis  (Various)
11.  Document on Gain Matching and Calibration  (WDMR)
12.  Future Plans (incl. contacts with ANU and LBL)  (BRF)
13.  Any other business
14.  Date of next meeting

W.N. Catford (Secretary)